could you explain how you got t1= 2s and t2= 6s?
1
Expert's answer
2020-12-02T09:44:45-0500

 Different answers in physics are usually a result of quadratic equations. For example, if you throw a body at 30 m/s vertically upward, and ask yourself "What time will it take to reach 2 meters above the ground?", the solution will be like this:


"h=v_0t-\frac{gt^2}{2},\\\space\\
\frac{gt^2}{2}-v_0t+h=0."

We transformed a physical equation into a mathematical equation with greater powers of t on the left. Substitute values:


"5t^2-30t+2=0."

This is a quadratic equation Replace numbers like this:


"5=a, \\b=-30,\\c=2."

The solution for t is


"t=\frac{-b\pm\sqrt{b^2-4ac}}{2a}."

With numbers plugged in, it is


"t=\frac{-(-30)\pm\sqrt{(-30)^2-4\cdot5\cdot2}}{2\cdot5}=0.07\text{ s or 5.9 s}."